If Anna and Jessica can clean house together in 6hrs and Anna can clean alone in 8hrs how long would it take Jessica to clean?
Found 2 solutions by stanbon, TimothyLamb:
Answer by stanbon(75887)

If Anna and Jessica can clean house together in 6hrs and Anna can clean alone in 8hrs how long would it take Jessica to clean?
----
Together rate = 1/6 job/hr
Anna rate = 1/8 job/hr
Jessica rate = 1/x job/hr
--------
Equation:
rate + rate = together rate
1/8 + 1/x = 1/6
----
Multiply thru by 24x to get:
3x + 24 = 4x
------
x = 24 hrs (time for Jessica to do the job alone)
============================
Cheers,
Stan H.
=======================

Answer by TimothyLamb(4379)

---
rates:
x = anna
y = jess
---
x + y = 1/6 job/hr
x = 1/8 job/hr
---
x + y = 1/6
1/8 + y = 1/6
y = 1/6 - 1/8
y = 8/48 - 6/48
y = 2/48
y = 1/24
---
time = 1/rate
---
1/y = jessica's time to clean the house = 24 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php